OpenStudy (anonymous):

evaluate:∫(x^2 dx)/(x^3+ 4)...

OpenStudy (anonymous):

Note that if you'll differentiate the denominator, with respect to x, you'll get 3x^2

OpenStudy (anonymous):

Substitute x^3 + 4 = t 3x^2 dx = dt => x^2dx = dt/3 ∫(x^2 dx)/(x^3+ 4) = (1/3)*∫dt/t ∫(x^2 dx)/(x^3+ 4) = ln (x^3 + 4)/3 + C

OpenStudy (anonymous):

thank u very much...